OpenStudy (anonymous):

how would I solve for 8^1/4 x (1/4)^x/2 = 16 ^3/4

OpenStudy (anonymous):

I think you have to use a log, but I have no clue. Anyone?

OpenStudy (anonymous):

I need to find the value for x I think

OpenStudy (anonymous):

and make bases the same

OpenStudy (anonymous):

What base would you use? I was never very good at log functions :(

OpenStudy (anonymous):

I would use 2

OpenStudy (anonymous):

but I dont know what to do after

OpenStudy (anonymous):

Ok, could you start by making everything in base 2?

OpenStudy (anonymous):

i did. and afer I got 2^3/4 x (1/2^2) ^ x/2 = 2^12/4

OpenStudy (anonymous):

Ok, now I'm as stuck as you are.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Sorry, try to grab some attention from others who are good at logs.

OpenStudy (anonymous):

ok thanks anyways

OpenStudy (campbell_st):

just write them all as powers of 2\[(2^3)^{\frac{1}{4}} \times (2^{-2})^{\frac{x}{2}} = (2^4)^{\frac{3}{4}}\] using index laws for power of a power and multiplication you get \[2^{\frac{3}{4} - x} = 2^3\] so all you need to do is solve the equation 3/4 - x = 3

OpenStudy (anonymous):

explain the index laws for power of a power and multiplication

OpenStudy (anonymous):

That's the part I don't remember :(

OpenStudy (anonymous):

okay I tried solving for what you gave me but couldn't get the right answer

OpenStudy (tyteen4a03):

@teic85 Power of a power: (a^b)^c = a^ (bc) Or a^b * a^c = a ^ (b+c)

OpenStudy (no.identity05):

 x=2.25........ try to check my answer gyus